XYZ CORP

Exercise

NYSE

Date

Price

Price

Close

Calls

OCT

85

16 3/4

101 11/16

OCT

90

12

101 11/16

OCT

95

7 5/8

101 11/16

Puts

OCT

85

1/8

101 11/16

OCT

90

3/8

101 11/16

OCT

95

13/16

101 11/16

  1. If you establish a long straddle using the options with an 85-exercise price, what is your dollar gain or loss if at expiration XYZ is still trading at 101 11/16?
  1. If you establish a long strap using the options with an 85-exercise price, what is your dollar gain or loss if at expiration XYZ is still trading at 101 11/16?
  1. If you establish a long strip using the options with an 85 exercise price, what is your dollar gain or loss if at expiration XYZ is still trading at 101 11/16?
Answer #1

Per share

1
=MAX(101-85,0)+MAX(85-101,0)-16.75-0.125=-0.875

2
=2*MAX(101-85,0)+MAX(85-101,0)-2*16.75-0.125=-1.625

3
=MAX(101-85,0)+2*MAX(85-101,0)-16.75-2*0.125=-1

Per contract

1.

=-87.5

2.

=-162.5

3.

=-100